# Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxies

### v1

@@ 6.3.8 Environmental dependence on relaxation timescales
We find a strong and significant difference between the two halo masses considered (two-sample KS-test statistic $=0.29$ and p-value $=1.5\times 10^{-6}$) showing a positive correlation between halo mass and the normalised relaxation time (in terms of $\bar{t}_{\rm{torque}}$). This is seen in both the overall shape of the distributions in Figure 13 and in the median relaxation times which correspond to $1.7t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (low-mass) and $2.8t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (intermediate/high-mass). Low-mass halos are dominated by short-duration normalised relaxation times, suggesting processes that preserve unstable misalignments (e.g. smooth accretion) are rare in this mass regime. On the other hand, intermediate/high-mass halos have a considerably flatter distribution and dominate the long-duration normalised relaxation timescales beyond $\gtrsim 4t_{\rm{relax}}/\bar{t}_{\rm{torque}}$. This behaviour becomes more pronounced when considering central galaxies with $M_{\rm{200}}>10^{13.5}$ M (not shown).
Although not shown, we also find a strong positive relationship between normalised relaxation times (in terms of $\bar{t}_{\rm{torque}}$) and stellar mass within $2r_{50}$. $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ increases from 1.6 to 2.5 to 3.7 for mass ranges of $M_{\rm{*}}<10^{10}$ M, $10^{10}$ M $<M_{\rm{*}}<10^{11}$ M, and $M_{\rm{*}}>10^{11}$ M, respectively. Again, these results are largely unchanged when considering only co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths.
Together, these results point toward halo cooling (from misaligned halos) being an important driver of long-duration misalignments in higher-mass systems in eagle. However, we stress that a more thorough analysis needs to be done to confirm the significance of accretion from misaligned halo cooling as a driver of misalignments within eagle and similar cosmological simulations.
@@ 6.4 Incidence of mergers with misalignment formation
In this section, we briefly explore the incidence of mergers with the formation of unstable misalignments. For this analysis, we consider galaxies with $M_{*}\gtrsim 10^{10}$ M in order to more accurately resolve minor mergers in the merger tree. This gives us a sample of 718 relaxations, with statistics given in Table 5. The behaviour of median $t_{\rm{relax}}$ values for ETGs and LTGs are comparable to those of our original sample, with a trend toward a larger and flatter distribution of $t_{\rm{relax}}/\bar{t}_{\rm{dyn}}$ and $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ as expected from our results in Section 6.3.8.
The incidence of mergers and the formation of unstable misalignments will be sensitive to the time window over which we associate these two events. We search for mergers with stellar mass ratios of $\mu_{\rm{*}}>0.1$ within windows of 0.4 Gyr, 0.6 Gyr, and 1.0 Gyr centred on the first snipshot to enter the unstable [$20^{\circ}-160^{\circ}$] regime. As galaxies tend to become unstably misaligned quickly, this snipshot overwhelmingly coincides with the first unstable misalignment [$30^{\circ}-150^{\circ}$]. Focusing primarily on the 0.6 Gyr window size, we aim to capture unstable misalignment formations that are explicitly associated with a direct merger origin (such as the first unstable misalignment in Galaxy B in Figure 4) and exclude unstable misalignment formations associated with an accretion event from an initial fly-by (such as Galaxy A in Figure 4). By setting this window to 0.6 Gyr we search for mergers within $\approx 2$ snipshots on either side of the kinematic instability. This gives us a small degree of flexibility with regard to the exact timing of the merger according to the existing merger trees, and gives the gas a short window within which to settle and form an unstable misaligned disc. As minor merger signatures are typically visible for $\sim 0.5$ Gyr, we reason that any unstable misalignments associated with mergers in our sample would likely be classified similarly in optical counterparts.
@@
Figure 14: Pie charts visualizing the number of unstable misalignments associated with major and minor mergers for ETGs, LTGs, and our total sample of galaxies, all with $M_{*}>10^{10}$ M. Overall, we find $\approx 16.7\%$ of our relaxation sample associated with mergers within a 0.6 Gyr window centred on the first unstable snipshot.
In Figure 14 we visualise the fraction of unstable misalignments associated with major and minor mergers for our 0.6 Gyr window according to our total sample and with the morphology before any misalignment (e.g. ETG $\rightarrow$ LTG relaxations are classified under ‘ETG’). In total, $\approx 16.7\%$ of our sample is associated with a merger origin, with major mergers ($\approx 11.7\%$) dominating this fraction over minor mergers ($\approx 5.0\%$) in both morphologies. Of these, unstable misalignments in ETGs show slightly higher merger-origin fractions ($\approx 20.6\%$) over LTGs ($\approx 12.8\%$) which may be due to ETGs being more susceptible to misalignments from stochastic external accretion owing to lower gasSF fractions. Changing the window size does not change the relative proportions of major and minor mergers nor the higher incidence in ETGs, but does affect the incidence of mergers with $\approx 10.7\%$ and $\approx 21.6\%$ for windows of size 0.4 Gyr and 1.0 Gyr, respectively. These results imply that major mergers are more effective at driving unstable misalignments than minor mergers, as the latter may lack sufficient gas to meaningfully change the angular momentum of the in-situ gas.
We conclude that galaxy mergers are not the primary driver behind stellar-gas misalignments in eagle galaxies. This once again points toward significant contributions of other processes that form unstable misalignments, which are in agreement with results from Casanueva et al. (2022). However, given that misaligned halos are often associated with galaxy encounters and mergers (e.g. Bett & Frenk, 2012), external processes may remain dominant in misalignment formation but these may only be weakly connected temporally.
@@ 6.5 Discussion and comparison to existing results
As explored in detail in Davis & Bureau (2016), assuming misalignments to be formed primarily through mergers alongside theoretical relaxation timescales (in the absence of smooth gas accretion), analytical models significantly overproduce the fraction of counter-rotating (relaxed) ETGs and under-produce the fraction of unstable misaligned (un-relaxed) ETGs (Davis & Bureau, 2016). While shorter gas depletion timescales can be invoked to reproduce observed distributions, such that a misaligned disc is depleted before it can relax, this solution is not favoured. This is because it requires both a significantly higher merger rate ($\sim 5-10$ times larger than predicted from $\Lambda$CDM) to supply frequent misaligned gas alongside a higher efficiency of processes to deplete it (e.g. increased AGN activity or star-formation efficiency) that far exceeds values measured for typical ETGs.
@@
An alternative solution suggests that there may be additional processes forming misalignments besides mergers. For instance, galaxies that are already gas-poor (such as red-sequence ETGs; Davis et al., 2022) have a lower angular momentum threshold needed to misalign the gas disc. This may make these systems more susceptible to becoming misaligned from other sources of misaligned accretion (e.g. Lagos et al., 2018; Khim et al., 2021; Casanueva et al., 2022). Indeed, recent results from semi-analytical models and simulations have shown the formation pathways for misalignments to be significantly more diverse than previously assumed in earlier observational studies. Using the galform semi-analytical model (Cole et al., 2000) and assuming mergers to be the only source of misalignments, Lagos et al. (2015) found misalignments in only $\sim 2-5\%$ of ETGs at $z=0$. However, when additionally considering isotropic smooth accretion from halo cooling, the fraction of misaligned ETGs increased to $\sim 46\%$ of ETGs bringing this value significantly closer to observational results (e.g. Davis et al., 2011; Bryant et al., 2019; Ristea et al., 2022). More recently, Khim et al. (2021) use the horizon-agn simulation to investigate the origins and timescales of misalignments for a sample of $\approx 27,900$ galaxies. They report four main formation paths for misalignments: mergers ($\approx 35\%$), gas stripping and interactions with the environment ($\approx 23\%$), interactions with other galaxies ($\approx 21\%$), and secular evolution such as filament accretion or misaligned halo cooling ($\approx 21\%$).
@@
Overall, our results tend to be in agreement with the presence of additional drivers of misalignments, rather than a dominant population of significantly longer-lived unstable misalignments in ETGs. From our results in Section 6.3.3, we find that the vast majority of unstable misalignments in ETGs relax within $\sim 6-8\bar{t}_{\mathrm{dyn}}$. These relaxations do not show substantially enhanced relaxation times compared to LTGs once timescales are normalised with respect to $\bar{t}_{\mathrm{torque}}$. While we do see a population of fairly long-duration unstable misalignments with $\sim 1-2$ Gyr (of comparable duration to van de Voort et al., 2015), and a population with enhanced relaxation timescales of $>3\bar{t}_{\mathrm{torque}}$ beyond theoretical timescales (Tohline et al., 1982; Lake & Norman, 1983), these are not dominant among our sample. Furthermore, the relatively low incidence of mergers ($\approx 16.7\%$) with the formation of unstable misalignments in ETGs suggests misalignments form through diverse formation pathways and may occur more frequently than previously assumed. Thus, at least in eagle, we conclude that a population of long-duration unstable misalignments formed through mergers in ETGs, as suggested by Davis & Bureau (2016), is not the main driver behind the misalignment distributions at $z\sim 0.1$.
@@
In qualitative terms, our relaxation timescales agree with those investigated in horizon-agn by Khim et al. (2021). Using a sample of misaligned ($>30^{\circ}$) galaxies and characterising relaxation timescales as a population decay, they find longer characteristic decay timescales in ETGs ($\approx 2.34$ Gyr) compared to LTGs ($\approx 0.75$ Gyr). This gives ETGs a $\approx 3.4$ times longer misalignment duration. By tracing the longevity of individual unstable misalignments in eagle, we tend to find on average shorter relaxation timescales overall. However, we also find longer relaxation times in ETGs ($\approx 0.51$ Gyr) compared to LTGs ($\approx 0.38$ Gyr). Besides underlying differences in approach (e.g. use of a decay timescale) and the different mass ranges considered, the longer relaxation times found in horizon-agn may be due to numerous reasons. Firstly, horizon-agn contains a larger number of $M_{*}>10^{10}$ M galaxies ($\approx 28,000$) compared to eagle ($\approx 3600$) at $z\sim 0$. These galaxies typically reside in denser environments in horizon-agn compared to eagle (e.g. Khim et al., 2020). This, in turn, may increase relaxation times due to increased galaxy encounters and smooth accretion events (Khim et al., 2021). Secondly, as the time intervals between successive simulation outputs is about twice as long in horizon-agn compared to eagle, galaxies in our sample are required to meet our sample criteria over a greater number of consecutive snapshots. The lower time resolution in horizon-agn may also obfuscate many shorter-duration misalignments that we are able to resolve. Thirdly, Khim et al. (2021) make use of projected 2D misalignment angles. While we found no substantial changes in the shape of the misalignment distribution at $z=0.1$ between 2D and 3D misalignment angles (see Figure 2), it is clear that the use of 3D misalignment angles is preferred when characterising relaxation timescales with respect to theoretical predictions (see Section 2). Finally and most importantly, while we aim to measure the relaxation timescales of unstable misalignments ($30^{\circ}-150^{\circ}$), Khim et al. (2021) measure the longevity of misaligned galaxies ($>30^{\circ}$). As a consequence, galaxies that spend considerable time in the stable counter-rotating regime (e.g. Galaxy B in Figure 4) would have characteristic misalignment timescales ($\approx 4.5$ Gyr) far exceeding the relaxation timescales that we measure ($\approx 0.35$ Gyr and $\approx 2.55$ Gyr).
@@
Bryant et al. (2019) highlight that it is difficult to achieve the observed dichotomy between ETG and LTG misalignment distributions at $z\approx 0.1$ through differences in theoretical relaxation timescales. Using a slightly modified expression for $\bar{t}_{\rm{torque}}$ and assuming a standard ellipticity of 0.2 and 0.8 for ETGs and LTGs respectively, they find a theoretically maximum 2.7×2.7$\times$2.7 × longer relaxation time in ETGs. Thus, if morphology is the primary driver of the observed dichotomy and assuming unstable misalignments are formed at the same rate in ETGs and LTGs, we should expect <2.7×<2.7$\times$< 2.7 × more misaligned ETGs than LTGs. Further considering variations in the typical size and radial velocity of gas discs and taking ellipticity values from a full realistic population of galaxies, this ratio drops to $\lesssim 1.4$. In eagle, we find a similarly low ratio of 1.3$\approx$ 1.3$\times$ 1.3 × (1.6$\approx$ 1.6$\times$ 1.6 × for co $\rightarrow$ co relaxation paths) between median relaxation times of ETGs and LTGs. Crucially, these ratios are significantly lower than the observed ratio of misaligned gas between ETGs and LTGs ($15\pm 7$) in SAMI and similar studies. Likewise, we find a misaligned gasSF ratio of $6.2\pm 1.2$ between ETGs and LTGs (see Figure 2). Thus, we concur with Bryant et al. (2019) that differences in theoretical relaxation timescales between ETGs and LTGs are not the dominant driver behind the observed misalignment distributions. Instead, we suggest that a higher frequency of misalignment events in ETGs is likely a dominant driver behind the observed misalignment distributions.
@@
We find a relatively low incidence of mergers and the formation of unstable misalignments ($\approx 11-22\%$), highlighting the presence of additional processes forming unstable misalignments. These results are also in qualitative agreement with recent work in cosmological simulations (e.g. Starkenburg et al., 2019; Khim et al., 2021; Casanueva et al., 2022), as well as semi-analytical models (e.g. Lagos et al., 2015). For instance, in earlier work using eagle, Casanueva et al. (2022) report only $\approx 4\%$ of misalignments ($>30^{\circ}$) coinciding with mergers and highlight the importance of gas depletion (possibly through outflows or stripping) to form misalignments through galaxy interactions and encounters. Quantitative differences are likely caused by our different methods (as we use eagle snipshots and utilise a time window within which to identify a merger-driven origin). The mass range of galaxies considered is also different, with Casanueva et al. (2022) including many dwarf galaxies ($M_{*}>10^{9}$ M) that may be located in isolated environments with few merger opportunities. Our results are also in good agreement with Khim et al. (2021) who found a major (minor) merger-incidence of $\sim 16\%$ ($\sim 15\%$) using a modified major and minor merger ratio definition of $\mu_{*}>0.25$ and $0.020.25>\mu_{*}>$ 0.02, respectively. For comparison, we find a major (minor) merger-incidence of $\approx 11.7\%$ ($\approx 5.0\%$). Quantitative differences likely arise from the inclusion of smaller merger ratios in their method alongside the aforementioned higher environment density of galaxies in horizon-agn, prompting more frequent mergers and galaxy encounters. Our results also show some agreement with recent observational results in SAMI, with a likely merger-origin associated with $\approx 14\%$ of misalignments in galaxies with a mass range of $10^{9}\lesssim M_{*}\lesssim 10^{11.5}$ M (Ristea et al., 2022).
@@
Finally, we find a qualitative agreement with Khim et al. (2021) on the relation between long-lived misalignments and environment density. While we find no significant difference between the $t_{\rm{relax}}$ of low and medium/high-mass halos (as a proxy for environment density), we do find a strong positive correlation between $t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ and halo mass. We also find a positive correlation between $t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ and higher ongoing gas accretion rates among central galaxies. Alongside Khim et al. (2021), we attribute the positive correlation between environment density and longer-lived unstable misalignments to both a higher incidence of interactions and gas stripping events in denser environments, as well as higher quantities of halo cooling. For the latter, an unstable misalignment in a more massive halo may be preserved through significant misaligned smooth gas accretion, provided the gaseous halo is misaligned. For instance, stellar-gas misalignments were found to be more common in galaxies with misaligned halos in the IllustrisTNG simulation (Duckworth et al., 2020a). The combination of lower gasSF masses and a higher frequency of halo angular momentum decoupling likely makes ETGs more prone to unstable misalignments from processes besides mergers. These conclusions are also in agreement with those by Casanueva et al. (2022).
@@
As highlighted in Appendix A, we find no substantial difference in the number of long-lived ($t_{\mathrm{relax}}>0.1$ Gyr) unstable misalignments identified through use of a 0, 0.1 or 0.2 Gyr latency time. Thus, we rule out the possibility of missing a significant number of long-lived unstable misalignments due to our choice of $t_{\mathrm{lat}}=0.1$ Gyr. Finally, our sample redshift window of $10<z<$ 1 ($\approx 8$ Gyr) naturally limits the number of complete $t_{\mathrm{relax}}\gtrsim 20t_{\mathrm{dyn}}$ misalignments we can measure from formation to relaxation, with $t_{\mathrm{relax}}\approx 40t_{\mathrm{dyn}}$ being an approximate upper limit. However, considering the limitations imposed by our relaxation sample criteria, the fact we still find a moderate population of long-lived misalignments suggests that misalignment-preserving processes are nevertheless present and that these may be more widespread than previously assumed.

### v2

@@ 6.3.8 Halo mass dependence on relaxation timescales
We find a strong and significant difference between the two halo masses considered (two-sample KS-test statistic =0.29=0.29 and p-value $1.5\times 10^{-6}$) showing a positive correlation between halo mass and the normalised relaxation time (in terms of $\bar{t}_{\rm{torque}}$). This is seen in both the overall shape of the distributions in Figure 14 and in the median relaxation times which correspond to $1.7t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (low-mass) and $2.8t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (intermediate/high-mass). Low-mass halos are dominated by short-duration normalised relaxation times, suggesting processes that preserve unstable misalignments (e.g. smooth accretion) are rare in this mass regime. On the other hand, intermediate/high-mass halos have a considerably flatter distribution and dominate the long-duration normalised relaxation timescales beyond $\gtrsim 4t_{\rm{relax}}/\bar{t}_{\rm{torque}}$. This behaviour becomes more pronounced when considering central galaxies with $M_{\rm{200c}}>10^{13.5}$ M (not shown).
Although not shown, we also find a strong positive relationship between normalised relaxation times (in terms of $\bar{t}_{\rm{torque}}$) and stellar mass within $2r_{50}$. $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ increases from 1.6 to 2.5 to 3.7 for mass ranges of $M_{\rm{*}}<10^{10}$ M, $10^{10}$ M $M_{\rm{*}}<10^{11}$ M, and $M_{\rm{*}}>10^{11}$ M, respectively. Again, these results are largely unchanged when considering only co $\rightarrow$ co and counter $\rightarrow$ counter relaxation paths.
Together, these results point toward halo cooling (from misaligned halos; see e.g. Hill et al., 2021) being an important driver of long-duration misalignments in higher-mass systems in eagle. However, we stress that a more thorough analysis needs to be done to confirm the significance of accretion from misaligned halo cooling as a driver of misalignments within eagle and similar cosmological simulations.
@@ 6.4 Incidence of mergers with misalignment formation
In this section, we briefly explore the incidence of mergers with the formation of unstable misalignments. For this analysis, we consider galaxies with $M_{*}\gtrsim 10^{10}$ M in order to more accurately resolve minor mergers in the merger tree. This gives us a sample of 718 relaxations, with statistics given in Table 5. The behaviour of median $t_{\rm{relax}}$ values for ETGs and LTGs are comparable to those of our original sample, with a trend toward a larger and flatter distribution of $t_{\rm{relax}}/\bar{t}_{\rm{dyn}}$ and $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ as expected from our results in Section 6.3.8. Alongside higher gas fractions and SFRs, galaxies at $z\approx 1$ experience more frequent galaxy interactions and mergers compared to $z\approx 0$ counterparts (Fensch et al., 2017; Pfeffer et al., 2023). As such, we further sub-divide our sample between high-$z$ ($0.35<z<1.00$) and low-$z$ ($z<0.35$) with sub-samples of 449 and 269 relaxations, respectively. These redshift ranges correspond to approximately equal temporal ranges of $\sim 4$ Gyr.
The incidence of mergers and the formation of unstable misalignments will be sensitive to the time window over which we associate these two events. We search for mergers with stellar mass ratios of $\mu_{\rm{*}}>0.1$ within windows of 0.4 Gyr, 0.6 Gyr, and 1.0 Gyr centred on the first snipshot to enter the unstable $20^{\circ}-160^{\circ}$] regime. As galaxies tend to become unstably misaligned quickly, this snipshot overwhelmingly coincides with the first unstable misalignment $30^{\circ}-150^{\circ}$]. Focusing primarily on the 0.6 Gyr window size, we aim to capture unstable misalignment formations that are explicitly associated with a direct merger origin (such as the first unstable misalignment in Galaxy B in Figure 4) and exclude unstable misalignment formations associated with an accretion event from an initial fly-by (such as Galaxy A in Figure 4). By setting this window to 0.6 Gyr we search for mergers within $\approx 2$ snipshots on either side of the kinematic instability. This gives us a small degree of flexibility with regard to the exact timing of the merger according to the existing merger trees, and gives the gas a short window within which to settle and form an unstable misaligned disc. Because minor merger signatures are typically visible for $\sim 0.5$ Gyr in observational studies (e.g. Lotz et al., 2010), we reason that any unstable misalignments associated with mergers in our sample would likely be classified similarly in optical counterparts.
@@
Figure 15: Pie charts visualizing the number of unstable misalignments associated with major and minor mergers for $0.35<z<1.00$ and $z<0.35$, with the latter further split between ETGs and LTGs, all with $M_{*}>10^{10}$ M. Overall, we find $\approx 10.4$ per cent ($\approx 20.5$ per cent) of our relaxation sample associated with mergers within redshift ranges $z<0.35$ ($0.35<z<1.00$).
In Figure 15 we visualise the fraction of unstable misalignments associated with major and minor mergers within a 0.6 Gyr window for sub-samples of high-$z$ and low-$z$ relaxing misalignments. For our low-$z$ sub-sample, we additionally show the merger dependence on the morphology before any misalignment (e.g. ETG $\rightarrow$ LTG relaxations are classified under ‘ETG’).
At low-$z$, $\approx 10.4$ per cent of our sample is associated with a merger origin, with major mergers ($\approx 7.1$ per cent) dominating this fraction over minor mergers ($\approx 3.3$ per cent) in both morphologies. Of these, unstable misalignments in ETGs show higher merger-origin fractions ($\approx 13.9$ per cent) over LTGs ($\approx 6.4$ per cent) which may be due to ETGs being more susceptible to misalignments from stochastic external accretion owing to lower gasSF fractions. Changing the window size does not change the relative proportions of major and minor mergers nor the higher incidence in ETGs, but does affect the incidence of mergers with $\approx 4.4$ per cent and $\approx 14.5$ per cent (both within $z<0.35$) for windows of size 0.4 Gyr and 1.0 Gyr, respectively. Among our high-$z$ sub-sample, we find a higher incidence of mergers $\approx 20.5$ per cent) though they remain sub-dominant in misalignment formation. Together with our low-$z$ sub-sample, we find an incidence of $\approx 16.7$ for relaxations in $0<z<1$. These results imply that major mergers are more effective at driving unstable misalignments than minor mergers, as the latter may lack sufficient gas to meaningfully change the angular momentum of the in-situ gas.
We conclude that galaxy mergers are not the primary driver behind stellar-gas misalignments in eagle galaxies, but they may have been more important at higher redshifts. This once again points toward significant contributions of other processes that form unstable misalignments, which are in agreement with results from Casanueva et al. (2022). However, given that misaligned halos are often associated with galaxy encounters and mergers (e.g. Bett & Frenk, 2012), external processes may remain dominant in misalignment formation but these may only be weakly connected temporally.
@@ 6.5 Discussion and comparison to existing results
[…]
@@
An alternative solution suggests that there may be additional processes forming misalignments besides mergers. For instance, galaxies that are already gas-poor (such as red-sequence ETGs; Davis et al., 2022) have a lower angular momentum threshold needed to misalign the gas disc. This may make these systems more susceptible to becoming misaligned from other sources of misaligned accretion (e.g. Lagos et al., 2018; Khim et al., 2021; Casanueva et al., 2022). Indeed, recent results from semi-analytical models and simulations have shown the formation pathways for misalignments to be significantly more diverse than previously assumed in earlier observational studies. Using the galform semi-analytical model (Cole et al., 2000) and assuming mergers to be the only source of misalignments, Lagos et al. (2015) found misalignments in only $\sim 2-5$ per cent of ETGs at $z=0$. However, when additionally considering isotropic smooth accretion from halo cooling, the fraction of misaligned ETGs increased to $\sim 46$ per cent of ETGs bringing this value significantly closer to observational results (e.g. Davis et al., 2011; Bryant et al., 2019; Ristea et al., 2022). More recently, Khim et al. (2021) use the horizon-agn simulation to investigate the origins and timescales of misalignments for a sample of $\approx 27,900$ galaxies. They report four main formation paths for misalignments: mergers ($\approx 35$ per cent), gas stripping and interactions with the environment ($\approx 23$ per cent), interactions with other galaxies ($\approx 21$ per cent), and secular evolution such as filament accretion or misaligned halo cooling ($\approx 21$ per cent).
@@
Overall, our results tend to be in agreement with the presence of additional drivers of misalignments, rather than a dominant population of significantly longer-lived unstable misalignments in ETGs. From our results in Section 6.3.3, we find that the vast majority of unstable misalignments in ETGs relax within $\sim 6-8\bar{t}_{\mathrm{dyn}}$. These relaxations do not show substantially enhanced relaxation times compared to LTGs once timescales are normalised with respect to $\bar{t}_{\mathrm{torque}}$. While we do see a population of fairly long-duration unstable misalignments with $\sim 1-2$ Gyr (of comparable duration to van de Voort et al., 2015), and a population with enhanced relaxation timescales of $3\bar{t}_{\mathrm{torque}}$ beyond theoretical timescales (Tohline et al., 1982; Lake & Norman, 1983), these are not dominant among our sample. Furthermore, the relatively low incidence of mergers ($\approx 10-21$ per cent for low-$z$ and high-$z$, respectively) with the formation of unstable misalignments in ETGs suggests misalignments form through diverse formation pathways and may occur more frequently than previously assumed. Thus, at least in eagle, we conclude that a population of long-duration unstable misalignments formed through mergers in ETGs, as suggested by Davis & Bureau (2016), is not the main driver behind the misalignment distributions at $z\sim 0.1$.
@@
In qualitative terms, our relaxation timescales agree with those investigated in horizon-agn by Khim et al. (2021). Using a sample of misaligned ($30^{\circ}$) galaxies and characterising relaxation timescales as a population decay, they find longer characteristic decay timescales in ETGs ($\approx 2.34$ Gyr) compared to LTGs ($\approx 0.75$ Gyr). This gives ETGs a $\approx 3.4$ times longer misalignment duration. By tracing the longevity of individual unstable misalignments in eagle, we tend to find on average shorter relaxation timescales overall. However, we also find longer relaxation times in ETGs ($\approx 0.51$ Gyr) compared to LTGs ($\approx 0.38$ Gyr). Besides underlying differences in approach (e.g. use of a decay timescale) and the different mass ranges considered, the longer relaxation times found in horizon-agn may be due to numerous reasons. Firstly, horizon-agn contains a larger number of $M_{*}>10^{10}$ M galaxies ($\approx 28,000$) compared to eagle ($\approx 3600$) at $z\sim 0$. These galaxies typically reside in denser environments in horizon-agn compared to eagle (e.g. Khim et al., 2020). This, in turn, may increase relaxation times due to increased galaxy encounters and smooth accretion events (Khim et al., 2021). Secondly, as the time intervals between successive simulation outputs is about twice as long in horizon-agn compared to eagle, galaxies in our sample are required to meet our sample criteria over a greater number of consecutive snapshots. The lower time resolution in horizon-agn may also obfuscate many shorter-duration misalignments that we are able to resolve. Thirdly, Khim et al. (2021) make use of projected 2D misalignment angles. While we found no substantial changes in the shape of the misalignment distribution at $z=0.1$ between 2D and 3D misalignment angles (see Figure 2), it is clear that the use of 3D misalignment angles is preferred when characterising relaxation timescales with respect to theoretical predictions (see Section 2). Finally and most importantly, while we aim to measure the relaxation timescales of unstable misalignments ($30^{\circ}-150^{\circ}$), Khim et al. (2021) measure the longevity of misaligned galaxies ($30^{\circ}$). As a consequence, galaxies that spend considerable time in the stable counter-rotating regime (e.g. Galaxy B in Figure 4) would have characteristic misalignment timescales ($\approx 4.5$ Gyr) far exceeding the relaxation timescales that we measure ($\approx 0.35$ Gyr and $\approx 2.55$ Gyr).
@@
Bryant et al. (2019) highlight that it is difficult to achieve the observed dichotomy between ETG and LTG misalignment distributions at $z\approx 0.1$ through differences in theoretical relaxation timescales. Using a slightly modified expression for $\bar{t}_{\rm{torque}}$ and assuming a standard ellipticity of 0.2 and 0.8 for ETGs and LTGs respectively, they find a theoretically maximum $2.7\times$ longer relaxation time in ETGs. Thus, if morphology is the primary driver of the observed dichotomy and assuming unstable misalignments are formed at the same rate in ETGs and LTGs, we should expect $2.7\times$ more misaligned ETGs than LTGs. Further considering variations in the typical size and radial velocity of gas discs and taking ellipticity values from a full realistic population of galaxies, this ratio drops to $\lesssim 1.4$. In eagle, we find a similarly low ratio of $\approx 1.3\times$ ($\approx 1.6\times$ for co $\rightarrow$ co relaxation paths) between median relaxation times of ETGs and LTGs. Crucially, these ratios are significantly lower than the observed ratio of misaligned gas between ETGs and LTGs ($15\pm 7$) in SAMI and similar studies. Likewise, we find a misaligned gasSF ratio of $6.2\pm 1.2$ between ETGs and LTGs (see Figure 2). Thus, we concur with Bryant et al. (2019) that differences in theoretical relaxation timescales between ETGs and LTGs are not the dominant driver behind the observed misalignment distributions. Instead, we suggest that a higher frequency of misalignment events in ETGs is likely the dominant driver behind the observed misalignment distributions.
@@
We find a relatively low incidence of mergers coinciding with the formation of unstable misalignments ($\approx 11-22$ per cent) over $0<z<1$, highlighting the presence of additional processes forming unstable misalignments, especially within $z<0.35$. These results are also in qualitative agreement with recent work in cosmological simulations (e.g. Starkenburg et al., 2019; Khim et al., 2021; Casanueva et al., 2022), as well as semi-analytical models (e.g. Lagos et al., 2015). For instance, in earlier work using eagle, Casanueva et al. (2022) report only $\approx 4$ per cent of misalignments ($30^{\circ}$) coinciding with mergers and highlight the importance of gas depletion (possibly through outflows or stripping) to form misalignments through galaxy interactions and encounters. Quantitative differences are likely caused by our different methods (as we use eagle snipshots and utilise a time window within which to identify a merger-driven origin). The mass range of galaxies considered is also different, with Casanueva et al. (2022) including many dwarf galaxies ($M_{*}>10^{9}$ M) that may be located in isolated environments with few merger opportunities.
Our results are also in relatively good qualitative agreement with Khim et al. (2021) who found a major (minor) merger-incidence of $\sim 16$ per cent ($\sim 15$ per cent) at $z\approx 0$ using a modified major and minor merger ratio definition of $\mu_{*}>0.25$ and $0.25>\mu_{*}>0.02$, respectively. For comparison, we find a major (minor) merger-incidence of $\approx 7.1$ per cent ($\approx 3.3$ per cent) for galaxies within $z<0.35$. Quantitative differences likely arise from the inclusion of smaller merger ratios in their method alongside the aforementioned higher environment density of galaxies in horizon-agn, prompting more frequent mergers and galaxy encounters. Our results also show some agreement with recent observational results in SAMI within $z<0.1$, with a likely merger-origin associated with $\approx 14$ per cent of misalignments in galaxies with a mass range of $10^{9}\lesssim M_{*}\lesssim 10^{11.5}$ M (Ristea et al., 2022).
@@
Finally, we find a qualitative agreement with Khim et al. (2021) on the relation between long-lived misalignments and halo mass as a proxy for environment density. While we find no significant difference between the $t_{\rm{relax}}$ of low and medium/high-mass halos, we do find a strong positive correlation between $t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ and halo mass. We also find a positive correlation between $t_{\mathrm{relax}}/\bar{t}_{\mathrm{torque}}$ and higher ongoing gas accretion rates among central galaxies. Alongside Khim et al. (2021), we attribute the positive correlation between halo mass and longer-lived unstable misalignments to both a higher incidence of interactions and gas stripping events in group environments, as well as higher quantities of halo cooling (and resulting gas inflow) with increasing halo mass (e.g. Correa et al., 2018; Mitchell & Schaye, 2022). For the latter, an unstable misalignment in a more massive halo may be preserved through significant misaligned smooth gas accretion, provided the gaseous halo is misaligned. For instance, stellar-gas misalignments were found to be more common in galaxies with misaligned halos in the IllustrisTNG simulation (Duckworth et al., 2020a). The combination of lower gasSF masses and a higher frequency of halo angular momentum decoupling likely makes ETGs more prone to unstable misalignments from processes besides mergers. These conclusions are also in agreement with those by Casanueva et al. (2022).
@@
As highlighted in Appendix A, we find no substantial difference in the number of long-lived ($t_{\mathrm{relax}}>0.1$ Gyr) unstable misalignments identified through use of a 0, 0.1 or 0.2 Gyr latency time. Thus, we rule out the possibility of missing a significant number of long-lived unstable misalignments due to our choice of $t_{\mathrm{lat}}=0.1$ Gyr. Finally, our sample redshift window of $0<z<1$ ($\approx 8$ Gyr) naturally limits the number of complete $t_{\mathrm{relax}}\gtrsim 20t_{\mathrm{dyn}}$ misalignments we can measure from formation to relaxation, with $t_{\mathrm{relax}}\approx 40t_{\mathrm{dyn}}$ being an approximate upper limit. However, considering the limitations imposed by our relaxation sample criteria, the fact we still find a moderate population of long-lived misalignments suggests that misalignment-preserving processes are nevertheless present and that these may be more widespread than previously assumed.